OpenStudy (anonymous):

What are the zeroes of the function? Graph the function y = (x + 3)(x – 4)(x –5).

OpenStudy (anonymous):

zeros: x= -3, 4, 5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

put these factors equal to zero (x + 3)(x – 4)(x –5)=0 now separately consider each factor equal to zero like x+3 =0 x-4=0 x-5 =0 now find each x